There is 95 pounds and 5 boys each boy gets get double the amount that a girl gets there are 7 girls how much does each girl get?
Answer by Diana541(24)

Here is the given information:
95 pounds
7 girls, 5 boys.
Each boy gets double.
So if each boy gets double, we can represent each boy as 2, and each girl as one. So:
7 girls
7*1=7
5 boys
5*2=10
So the total given out is 17.
95/17= 5.5882 shows how much one is.
Since each girl only got one, then they got 5.5882 pounds.